# Building symmetries into data-driven manifold dynamics models for complex flows

### v1

@@ Abstract
Symmetries in a dynamical system provide an opportunity to dramatically improve the performance of data-driven models. For fluid flows, such models are needed for tasks related to design, understanding, prediction, and control. In this work we exploit the symmetries of the Navier-Stokes equations (NSE) and use simulation data to find the manifold where the long-time dynamics live, which has many fewer degrees of freedom than the full state representation, and the evolution equation for the dynamics on that manifold. We call this method “symmetry charting”. The first step is to map to a “fundamental chart”, which is a region in the state space of the flow to which all other regions can be mapped by a symmetry operation. To map to the fundamental chart we identify a set of indicators from the Fourier transform that uniquely identify the symmetries of the system. We then find a low-dimensional coordinate representation of the data in the fundamental chart with the use of an autoencoder. We use a variation called an implicit rank minimizing autoencoder with weight decay, which in addition to compressing the dimension of the data, also gives estimates of how many dimensions are needed to represent the data: i.e. the dimension of the invariant manifold of the long-time dynamics. Finally, we learn dynamics on this manifold with the use of neural ordinary differential equations. We apply symmetry charting to two-dimensional Kolmogorov flow in a chaotic bursting regime. This system has a continuous translation symmetry, and discrete rotation and shift-reflect symmetries. With this framework we observe that less data is needed to learn accurate data-driven models, more robust estimates of the manifold dimension are obtained, equivariance of the NSE is satisfied, better short-time tracking with respect to the true data is observed, and long-time statistics are correctly captured.
@@ I Introduction
In recent years, neural networks (NNs) have been implemented to learn data-driven low-dimensional representations and dynamical models of flow problems, with success in systems including the Moehlis-Faisst-Eckhardt (MFE) model [1], Kolmogorov flow [2, 3], and minimal turbulent channel flow [4]. For the most part however, these approaches do not explicitly take advantage of the fact that for dissipative systems like the Navier-Stokes Equations (NSE), the long-time dynamics are expected to lie on an invariant manifold $\mathcal{M}$ (sometimes called an inertial manifold [5, 6, 7]), $d_{\mathcal{M}}$, whose dimension may be much smaller than the nominal number of degrees of freedom $N$ required to specify the state of the system. Ideally, one could identify $d_{\mathcal{M}}$ from data, find a coordinate representation for points on $\mathcal{M}$, and learn the time-evolution on $\mathcal{M}$ in those coordinates. This would be a minimal-dimensional data-driven dynamic model. Linot & Graham explored this approach for chaotic dynamics of the Kuramoto-Sivashinsky equation (KSE) [8, 9]. They showed that the mean-squared error (MSE) of the reconstruction of the snapshots using an autoencoder (AE) for dimension reduction for the domain size of $L=22$ exhibited an orders-of-magnitude drop when the dimension of the inertial manifold is reached. Furthermore, modeling the dynamics with a dense NN at this dimension either with a discrete-time map [8] or a system of ordinary differential equations (ODE) [9] yielded excellent trajectory predictions and long-time statistics. The approach they introduced is referred to here as data-driven manifold dynamics (DManD). For the KSE in larger domains, it was found that simply observing MSE vs. autoencoder bottleneck dimension was not sufficient to determine the manifold dimension and exhaustive tests involving time evolution models vs. dimension were necessary to estimate the manifold dimension. Pérez De Jesús & Graham extended this approach to two-dimensional Kolmogorov flow in a chaotic regime [10]. Here dimension reduction from 1024 dimensions to $\leq 10$ was achieved, with very good short-time tracking predictions, long-time statistics, as well as accurate predictions of bursting events. Linot & Graham considered three-dimensional direct numerical simulations of turbulent Couette flow at $\text{Re}=400$ and found accurate data-driven dynamic models with fewer than 20 degrees of freedom [11]. These models were able to capture characteristics of the flows such as streak breakdown and regeneration, short-time tracking, as well as Reynolds stresses and energy balance. They also computed unstable periodic orbits from the models with close resemblance to previously computed orbits from the full system. Relatedly, Zeng et al. [12] exploited advances in autoencoder architecture [13] to yield more precise estimates of $d_{\mathcal{M}}$ for data from high-dimensional chaotic systems. The present work illustrates how symmetries of a flow system can be exploited in the DManD framework to yield highly efficient low-dimensional data-driven models for chaotic flows.
@@
A fundamental notion in the topology of manifolds, which will be useful in exploiting symmetry, is that of charts and atlases [14]. Simply put, a chart is a region of a manifold whose points can be represented in a local Cartesian coordinate system of dimension $d_{\mathcal{M}}$ and which overlaps with neighboring charts, while an atlas is a collection of charts that covers the manifold. This representation of a manifold has several advantages. First, for a manifold with dimension $d_{\mathcal{M}}$, it may not be possible to find a global coordinate representation in $d_{\mathcal{M}}$ dimensions: Whitney’s embedding theorem states that generic smooth maps for a smooth manifold of dimension $d_{\mathcal{M}}$ can be embedded into a Euclidean space of $2d_{\mathcal{M}}$. Dividing a manifold into an atlas of charts enables minimal-dimensional representations locally in $d_{\mathcal{M}}$ dimensions. Second, from the dynamical point of view, dynamics on different parts of a manifold may be very different and a single global representation of a manifold and the dynamics on it may not efficiently capture the dynamics, especially in the data-driven context. These two advantages of an atlas-of-charts representation have been exploited for data-driven modeling. Floryan & Graham developed a method to implement data-driven local representations for dynamical systems such as the quasiperiodic dynamics on a torus, a reaction-diffusion system, and the KSE to learn dynamics on invariant manifolds of minimal dimension [15]. They refer to this method as Charts and Atlases for Nonlinear Data-Driven Dynamics on Manifolds – “CANDyMan”. Fox et al. then applied this to the MFE model, which displays highly intermittent behavior in the form of quasilaminarization and full relaminarization events, demonstrating more accurate time evolution predictions as global (“single chart”) model [16]. In the present work, we use the charts and atlases framework to exploit symmetries.
@@
In the Navier-Stokes equations (NSE), symmetries appear in the form of continuous and discrete symmetry groups. The symmetries of the NSE in physical space are reflected in state space as symmetries of the vector field, and these can generate natural charts of the system. Consider the vector field $\boldsymbol{q}$ shown in Fig. 1a where the operation $\mathcal{R}\boldsymbol{q}$ rotates the vector field by $\pi/2$ in the clockwise direction. Trajectories of the ODE $\dot{\boldsymbol{y}}=\boldsymbol{q}$ will reflect this symmetry. Without exploiting symmetry, any model trained on this data will need to represent the whole state space and even so will generally not obey the exact symmetries of the true system due to finite sampling effects. In this work we leverage knowledge of the charts by explicitly considering and factoring out the symmetries. We show a depiction of this in Fig. 1b where we know that by applying $\mathcal{R}$ to the data we can map it to a different quadrant, and that by mapping points in quadrants 2, 3, and 4 with $\mathcal{R}$, $\mathcal{R}^{2}$, and $\mathcal{R}^{3}$, respectively, they all end up in quadrant 1. This is shown in Fig. 1c where the data collapses on top of each other when mapped to the first quadrant. We will call this region the “fundamental domain” for the state space, and when expanded to overlap with its neighbors, the “fundamental chart”.
@@
Figure 1: (a) Vector field $\boldsymbol{q}$ with symmetry group $\mathcal{R}$ in state space. (b) Data in each quadrant can be mapped to another quadrant by operating with $\mathcal{R}$. (c) After mapping everything to the positive quadrant data lies on top of each other.
In the present work, we present a framework for learning DManD models in the fundamental chart, as in Fig. 1c, where the dynamics of the flow occur, and apply it to chaotic dynamics of the NSE. We refer to this method as “symmetry charting”, which results in (at least approximately) minimal-dimensional data-driven models that are equivariant to symmetry operations. In addition the model only need to be learned in one chart because a combination of its symmetric versions will cover the manifold. We first learn minimal-dimensional representations in the fundamental chart followed by an evolution equation for the dynamics on it. To this end we use information about the symmetries of the system combined with NNs to learn the compressed representation and dynamics.
We apply symmetry charting to two-dimensional Kolmogorov flow in a regime where the chaotic dynamics travels between unstable relative periodic orbits (RPOs) [21] through bursting events [22] that shadow heteroclinic orbits connecting the RPOs. RPOs correspond to periodic orbits in a moving reference frame, such that in a fixed frame, the pattern at time $t+T$ is a phase-shifted replica of the pattern at time $t$. To achieve this goal we map the data to a fundamental chart which is then used to train undercomplete AEs to find low-dimensional representations. To this end, we use Implicit Rank Minimizing autoencoders with weight decay (IRMAE-WD) [12]. This architecture was introduced by Zeng & Graham as a method that estimates $d_{\mathcal{M}}$ without the need to sweep over dimensions and using MSE and statistics as indicators [8, 10]. We give an overview of IRMAE-WD in Section III.2. After learning a low-dimensional representation with IRMAE-WD we learn a time map with the use of the neural ODE (NODE) [23], which we discuss in Section III.3. We then show results in Section IV and finish with concluding remarks in Section V.
@@ II Kolmogorov flow, symmetries, and projections
The two-dimensional NSE with Kolmogorov forcing are
@@
where flow is in the $x-y$ plane, $\boldsymbol{u}=[u,v]$ is the velocity vector, $p$ is the pressure, $n$ is the wavenumber of the forcing, and $\hat{\boldsymbol{x}}$ is the unit vector in the $x$ direction. Here $\mathrm{Re}=\frac{\sqrt{\chi}}{v}\left(\frac{L_{y}}{2\pi}\right)^{3/2}$ where $\chi$ is the dimensional forcing amplitude, $\nu$ is the kinematic viscosity, and $L_{y}$ is the size of the domain in the $y$ direction. We consider the periodic domain $[0,2\pi/\alpha]\times[0,2\pi]$ with $\alpha=1$. Vorticity is defined as $\omega=\hat{\boldsymbol{z}}\cdot\boldsymbol{\nabla}\times\boldsymbol{u}$, where $\hat{\boldsymbol{z}}$ is the unit vector in the $z$ direction (orthogonal to the flow). The equations are invariant under several symmetry operations [24], namely a shift (in $y$)-reflect (in $x$), a rotation through $\pi$, and a continuous translation in $x$:
@@
For $n$ wavelengths the state can be shift-reflected $12n-$ 1 times with the operation $\mathscr{S}$. Together with the rotation $\mathscr{R}$ there will be a total of $4n$ states that are related by symmetries. In the case of $n=2$ this results in eight regions of state space that are related to one another by symmetries. In the chaotic regime that we consider, trajectories visit each of these regions. Relatedly, velocity fields for the Kolmogorov flow satisfy the equivarance property associated with these actions: if $\boldsymbol{u}$ is a solution at a given time, then so are $\mathscr{S}\boldsymbol{u}$, $\mathscr{R}\boldsymbol{u}$, and $\mathscr{T}_{l}\boldsymbol{u}$. We elaborate on this property in Sec. III.1.
@@
where $\left\langle\cdot\right\rangle_{V}$ corresponds to the average taken over the domain. For the case of $n=1$ the laminar state is linearly stable at all Re [25]. It is not until $n=2$ that the laminar state becomes unstable, with a critical value of $\mathrm{Re}_{c}=n^{3/2}2^{1/4}$[26, 27, 28]. In what follows, we evolve the NSE numerically in the vorticity representation on a $[d_{x}\times d_{y}]=[32\times 32]$ grid following the pseudo-spectral scheme given by Chandler & Kerswell [24], which is based on the code by Bartello & Warn [29].

### v2

@@ Abstract
Data-driven reduced-order models of the dynamics of complex flows are important for tasks related to design, understanding, prediction, and control. Many flows obey symmetries, and the present work illustrates how these can be exploited to yield highly efficient low-dimensional data-driven models for chaotic flows. In particular, incorporating symmetries both guarantees that the reduced order model automatically respects them and dramatically increases the effective density of data sampling. Given data for the long-time dynamics of a system, and knowing the set of continuous and discrete symmetries it obeys, the first step in the methodology is to identify a “fundamental chart”, a region in the state space of the flow to which all other regions can be mapped by a symmetry operation, and a set of criteria indicating what mapping takes each point in state space into that chart. We then find a low-dimensional coordinate representation of the data in the fundamental chart with the use of an autoencoder architecture that also provides an estimate of the dimension of the invariant manifold where data lie. Finally, we learn dynamics on this manifold with the use of neural ordinary differential equations. We apply this method, denoted “symmetry charting” to simulation data from two-dimensional Kolmogorov flow in a chaotic bursting regime. This system has a continuous translation symmetry, and discrete rotation and shift-reflect symmetries. With this framework we observe that less data is needed to learn accurate data-driven models, more robust estimates of the manifold dimension are obtained, equivariance of the NSE is satisfied, better short-time tracking with respect to the true data is observed, and long-time statistics are correctly captured.
@@ I Introduction
In recent years, neural networks (NNs) have been implemented to learn data-driven low-dimensional representations and dynamical models of flow problems, with success in systems including the Moehlis-Faisst-Eckhardt (MFE) model [1], Kolmogorov flow [2, 3], and minimal turbulent channel flow [4]. For the most part however, these approaches do not explicitly take advantage of the fact that for dissipative systems like the Navier-Stokes Equations (NSE), the long-time dynamics are expected to lie on an invariant manifold $\mathcal{M}$ (sometimes called an inertial manifold [5, 6, 7]), whose dimension $d_{\mathcal{M}}$ may be much smaller than the nominal number of degrees of freedom $N$ required to specify the state of the system. Ideally, one could identify $d_{\mathcal{M}}$ from data, find a coordinate representation for points on $\mathcal{M}$, and learn the time-evolution on $\mathcal{M}$ in those coordinates. This would be a minimal-dimensional data-driven dynamic model. Linot & Graham explored this approach for chaotic dynamics of the Kuramoto-Sivashinsky equation (KSE) [8, 9]. They showed that the mean-squared error (MSE) of the reconstruction of the snapshots using an autoencoder (AE) for dimension reduction for the domain size of $L=22$ exhibited an orders-of-magnitude drop when the dimension of the inertial manifold is reached. Furthermore, modeling the dynamics with a dense NN at this dimension either with a discrete-time map [8] or a system of ordinary differential equations (ODE) [9] yielded excellent trajectory predictions and long-time statistics. The approach they introduced is referred to here as data-driven manifold dynamics (DManD). For the KSE in larger domains, it was found that simply observing MSE vs. autoencoder bottleneck dimension was not sufficient to determine the manifold dimension and exhaustive tests involving time evolution models vs. dimension were necessary to estimate the manifold dimension. Pérez De Jesús & Graham extended this approach to two-dimensional Kolmogorov flow in a chaotic regime [10]. Here dimension reduction from 1024 dimensions to $\leq 10$ was achieved, with very good short-time tracking predictions, long-time statistics, as well as accurate predictions of bursting events. Linot & Graham considered three-dimensional direct numerical simulations of turbulent Couette flow at $\text{Re}=400$ and found accurate data-driven dynamic models with fewer than 20 degrees of freedom [11]. These models were able to capture characteristics of the flows such as streak breakdown and regeneration, short-time tracking, as well as Reynolds stresses and energy balance. They also computed unstable periodic orbits from the models with close resemblance to previously computed orbits from the full system. Relatedly, Zeng et al. [12] exploited advances in autoencoder architecture [13] to yield more precise estimates of $d_{\mathcal{M}}$ for data from high-dimensional chaotic systems. The present work illustrates how symmetries of a flow system can be exploited in the DManD framework to yield highly efficient low-dimensional data-driven models for chaotic flows. In particular, incorporating symmetries both guarantees that the reduced order model automatically respects them and dramatically increases the effective density of data sampling.
@@
A fundamental notion in the topology of manifolds, which will be useful in exploiting symmetry, is that of charts and atlases [14]. Simply put, a chart is a region of a manifold whose points can be represented in a local Cartesian coordinate system of dimension $d_{\mathcal{M}}$ and which overlaps with neighboring charts, while an atlas is a collection of charts that covers the manifold. (Strictly speaking, a chart is the region and the local coordinate representation of each point in the region. We do not emphasize the latter aspect here because we will always have at hand explicit coordinates.) This representation of a manifold has several advantages. First, for a manifold with dimension $d_{\mathcal{M}}$, it may not be possible to find a global coordinate representation in $d_{\mathcal{M}}$ dimensions: Whitney’s embedding theorem states that generic smooth maps for a smooth manifold of dimension $d_{\mathcal{M}}$ can be embedded into a Euclidean space of $2d_{\mathcal{M}}$. Dividing a manifold into an atlas of charts enables minimal-dimensional representations locally in $d_{\mathcal{M}}$ dimensions. Second, from the dynamical point of view, dynamics on different parts of a manifold may be very different and a single global representation of a manifold and the dynamics on it may not efficiently capture the dynamics, especially in the data-driven context. These two advantages of an atlas-of-charts representation have been exploited for data-driven modeling. Floryan & Graham developed a method to implement data-driven local representations for dynamical systems such as the quasiperiodic dynamics on a torus, a reaction-diffusion system, and the KSE to learn dynamics on invariant manifolds of minimal dimension [15]. They refer to this method as Charts and Atlases for Nonlinear Data-Driven Dynamics on Manifolds – “CANDyMan”. Fox et al. then applied this to the MFE model, which displays highly intermittent behavior in the form of quasilaminarization and full relaminarization events, demonstrating more accurate time evolution predictions as global (“single chart”) model [16]. In the present work, we use the charts and atlases framework to exploit symmetries.
In partial differential equations (PDEs) such as the Navier-Stokes equations (NSE), symmetries appear in the form of continuous and discrete operations such as translations, reflections, and rotations, under which the equations are unchanged, and which map solutions to solutions. For example, for flow in a duct with a rectangular cross-section, the reflection of a solution to the problem across the symmetry planes of the cross section is also a solution to the problem. For flow in a circular pipe, the rotation of a solution by any angle around the central axis is another solution, and likewise for reflection across any plane cutting through that axis. For Kolmogorov flow, the example considered here, the symmetries arise from the nature of the forcing and the periodic boundary conditions, as detailed in Section II. We emphasize that the approach described here is data-driven, relying only on knowledge the appropriate symmetries, not on the specific governing equations.
@@
Symmetries in physical space are reflected in state space as symmetries of the vector field, and these can generate natural charts of the system. In this work we leverage knowledge of the symmetries to construct reduced-order models that automatically satisfy them. As an illustrative example, consider points $\boldsymbol{y}$ in state space $\mathbb{R}^{2}$ (or equivalently, in a two-dimensional invariant manifold in a higher-dimensional state space) and the vector field $\boldsymbol{q}(\boldsymbol{y})$ shown in Fig. 1a. This field is equivariant with respect to the operation $\mathcal{R}\boldsymbol{q}$ that rotates the vector field by $\pi/2$ in the clockwise direction around the origin. That is, $\boldsymbol{q}(\mathcal{R}\boldsymbol{y})=\mathcal{R}\boldsymbol{q}(\boldsymbol{y})$. Trajectories of the ODE $\dot{\boldsymbol{y}}=\boldsymbol{q}$ will reflect this symmetry. Without exploiting symmetry, any model trained on trajectory data $\boldsymbol{y}(t)$ will need to represent the whole state space and even so will generally not obey the exact symmetries of the true system due to finite sampling effects. We show a depiction of in Fig. 1b where we know that by applying $\mathcal{R}$ to the data we can map it to a different quadrant, and that by mapping points $\boldsymbol{y}$ and corresponding vectors $\boldsymbol{q}$ in quadrants 2, 3, and 4 with $\mathcal{R}$, $\mathcal{R}^{2}$, and $\mathcal{R}^{3}$, respectively, they all end up in quadrant 1. This is shown in Fig. 1c, where the vectors collapse on top of each other when mapped to the first quadrant. We will call this region the “fundamental domain” for the state space (or an invariant manifold in the state space), and when expanded to overlap with its neighbors, the “fundamental chart” [17]. For each quadrant $j$ we can now define an “indicator” $\mathcal{I}=j-1$, such that a point $\boldsymbol{y}_{j}$ in quadrant $j=\mathcal{I}+1$ can be uniquely written in terms of a point $\boldsymbol{y}_{1}\equiv\boldsymbol{y}_{F}$ in the first quadrant – i.e. in the fundamental domain, with symmetries “factored out” – and the indicator $\mathcal{I}$, via the relation $\boldsymbol{y}_{F}=\mathcal{O}_{\mathcal{I}}\boldsymbol{y}_{\mathcal{I}+1}$, with $\mathcal{O}_{\mathcal{I}}=\mathcal{R}^{\mathcal{I}}$. Likewise for the corresponding vector field $\boldsymbol{q}$. That is, the “state-indicator” pair $\left\{\boldsymbol{y}_{F},\mathcal{I}\right\}$ uniquely determines each point in state space. Thus a trajectory $\boldsymbol{y}(t)$ that evolves under $\dot{\boldsymbol{y}}=\boldsymbol{q}$ can be written as $\left\{\boldsymbol{y}_{F}(t),\mathcal{I}(t)\right\}$. The present work generalizes this formalism.
@@
Figure 1: (a) Vector field $\boldsymbol{q}$ with discrete symmetry group $\mathcal{R}$ in state space or on an invariant manifold within a state space. (b) Points and vectors in each quadrant can be mapped to another quadrant by operating with $\mathcal{R}$. (c) Since the vector field at any point in the domain can be mapped into the first quadrant, this region of state space can be taken to be the fundamental domain. Extending its boundaries to overlap with the other quadrants yields the fundamentl chart for this vector field.
In the present work, we present a framework for learning DManD models in the fundamental chart, as in Fig. 1c, where the dynamics of the flow occur, and apply it to chaotic dynamics of the NSE. We refer to this method as “symmetry charting”, which results in (at least approximately) minimal-dimensional data-driven models that are equivariant to symmetry operations. In addition the model only need to be learned in one chart because a combination of its symmetric versions will cover the manifold. This approach is similar to the CANDyMan methodology, where now the charts are related by symmetry operations. We first learn minimal-dimensional representations in the fundamental chart followed by an evolution equation for the dynamics on it. To this end we use information about the symmetries of the system combined with NNs to learn the compressed representation and dynamics.
We apply symmetry charting to data from numerical simulations of two-dimensional Kolmogorov flow. This flow is chosen because it is well-studied as a model of a flow with complex dynamics [21, 22, 23, 24, 25, 26, 27, 2, 2, 10] while being relatively inexpensive to simulate. Additionally, an experimental realization of this flow has been developed [28]. We focus on a Reynolds number regime where chaotic trajectories travel between unstable relative periodic orbits (RPOs) [29] through bursting events [30] that shadow heteroclinic orbits connecting the RPOs. RPOs correspond to periodic orbits in a moving reference frame, such that in a fixed frame, the pattern at time $t+T$ is a phase-shifted replica of the pattern at time $t$. To perform symmetry charting we map the data to a fundamental chart which is then used to train undercomplete AEs to find low-dimensional representations. To this end, we use Implicit Rank Minimizing autoencoders with weight decay (IRMAE-WD) [13, 12]. This architecture enables estimation of $d_{\mathcal{M}}$ without the need to sweep over dimensions and using MSE and statistics as indicators [8, 10]. We give an overview of IRMAE-WD in Section III.2. After learning a low-dimensional representation with IRMAE-WD we learn a time map with the use of the neural ODE (NODE) technique[31], which we discuss in Section III.3. We then show results in Section IV and finish with concluding remarks in Section V. Appendix A includes details of applying symmetry operations in Fourier space, and Appendix B outlines how the symmetry charting method described here applies to 3D plane channel flow.
@@ II Kolmogorov flow, symmetries, and projections
The “data set” we consider in applying symmetry charting consists of snapshots of the vorticity field from numerical solutions of the two-dimensional (2D) Navier-Stokes equations with Kolmogorov forcing:
@@
where flow is in the $x-y$ plane, $\boldsymbol{u}=[u,v]$ is the velocity vector, $p$ is the pressure, $n$ is the wavenumber of the forcing, and $\hat{\boldsymbol{x}}$ is the unit vector in the $x$ direction. Here $\mathrm{Re}=\frac{\sqrt{\chi}}{v}\left(\frac{L_{y}}{2\pi}\right)^{3/2}$ where $\chi$ is the dimensional forcing amplitude, $\nu$ is the kinematic viscosity, and $L_{y}$ is the size of the domain in the $y$ direction. We consider the square periodic domain $[-\pi,\pi]\times[-\pi,\pi]$. Vorticity is defined as $\omega=\hat{\boldsymbol{z}}\cdot\boldsymbol{\nabla}\times\boldsymbol{u}$, where $\hat{\boldsymbol{z}}$ is the unit vector in the $z$ direction (orthogonal to the flow). Because of the symmetries of the forcing term $\sin(ny)\hat{\boldsymbol{x}}$, along with the rectangular periodic shape of the domain, the equations are invariant under several symmetry operations [22], namely a shift (in $y$)-reflect (in $x$) $\mathscr{S}$, a rotation $\mathscr{R}$ through $\pi$, and a continuous translation $\mathscr{T}_{l}$ over an arbitrary distance $l$ in $x$:
@@
For a forcing with $y$-wavenumber $n$, the state can be shift-reflected $12n-$ 1 times with the operation $\mathscr{S}$. Together with the rotation $\mathscr{R}$ there will be a total of $4n$ states that are related by symmetries. In the case of $n=2$ this results in eight regions of state space that are related to one another by these discrete symmetries. In the chaotic regime that we consider, trajectories visit each of these regions. Relatedly, velocity fields for the Kolmogorov flow satisfy the equivarance property associated with these actions: if $\boldsymbol{u}$ is a solution at a given time, then so are $\mathscr{S}\boldsymbol{u}$, $\mathscr{R}\boldsymbol{u}$, and $\mathscr{T}_{l}\boldsymbol{u}$. We elaborate on this property in Sec. III.1.
@@
where $\left\langle\cdot\right\rangle_{V}$ corresponds to the average taken over the domain. For the case of $n=1$ the laminar state is linearly stable at all Re [32]. It is not until $n=2$ that the laminar state becomes unstable, with a critical value of $\mathrm{Re}_{c}=n^{3/2}2^{1/4}$[21, 33, 25]. In what follows, we evolve the NSE numerically in the vorticity representation on a $[d_{x}\times d_{y}]=[32\times 32]$ grid following the pseudo-spectral scheme given by Chandler & Kerswell [22], which is based on the code by Bartello & Warn [34]. We find that this resolution is sufficient at the conditions considered. We emphasize however, that the aim of the present work is not to use machine learning to solve the NSE, but rather to find a high-fidelity low-dimensional representation of the dynamics from a data set, which in this case comes from the numerical solution described here.